OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (anonymous):

@TuringTest @TheSmartOne @King.Void.

OpenStudy (king.void.):

Not good at Math. Sorry.

OpenStudy (anonymous):

ok thank you anyways @King.Void.

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi!!!

OpenStudy (misty1212):

it would be good to use the equation tool, so you don't have to say things like "three halves" that is why people invented numerals

OpenStudy (misty1212):

\[y=\frac{3}{2}x+4\]

OpenStudy (misty1212):

you know what the slope of that is?

OpenStudy (anonymous):

no @misty1212

OpenStudy (misty1212):

i bet you do! the slope of \(y=mx+b\) is \(m\) so the slope of \(y=\frac{3}{2}x+4\) is ?

OpenStudy (anonymous):

uhhh 3/2

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

srry math is not my greatest subject so please bear w. me

OpenStudy (misty1212):

next question is, what is the slope of the a line perpendicular to a line with slope \(\frac{3}{2}\) ?

OpenStudy (anonymous):

i have no clue may you explain to me more

OpenStudy (misty1212):

take the number \(\frac{3}{2}\) and flip it then put a minus sign in front it is called the "negative reciprocal"

OpenStudy (anonymous):

ok so th negative reciprocal is -2/3

OpenStudy (misty1212):

right that narrows your choices

OpenStudy (anonymous):



OpenStudy (anonymous):

ok so we will eliminate c and d right

OpenStudy (misty1212):

yea so it is one of the last two choices \[y=-\frac{2}{3}x+3\] or \[y=-\frac{2}{3}x-1\]

OpenStudy (misty1212):

because both of those have slope \(-\frac{2}{3}\) as we now know that the slope of \(y=mx+b\) is \(m\)

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

how are you going to choose?

OpenStudy (anonymous):

ummm find the slope-intercept form

OpenStudy (anonymous):

y= mx +b i think

OpenStudy (misty1212):

you can do that start with the point slope form \[y-3=-\frac{2}{3}(x+6)\] and solve for \(y\)

OpenStudy (anonymous):

ummm may you help me w. this

OpenStudy (anonymous):

y-3 =-2/3x -4 right

OpenStudy (misty1212):

yea, then add 3 to get \(y\) by itself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-3+3=-2/3x -4 + 3

OpenStudy (misty1212):

and...

OpenStudy (anonymous):

and???

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

what is \(-3+3\) ?

OpenStudy (anonymous):

0

OpenStudy (misty1212):

and what is \(-4+3\)?

OpenStudy (anonymous):

-1 @misty1212

OpenStudy (misty1212):

ok good then final answer is \[y=-\frac{2}{3}x-1\]

OpenStudy (anonymous):

that means my answer is d right

OpenStudy (anonymous):

thank you may you help me w .another and i will repost